# Quantifying the Impact of Incompleteness on Identifying and Interpreting Galaxy Protocluster Populations with the TNG-Cluster Simulation

### v1

@@ 2.3 Simulated Protocluster Population
In the review paper by Overzier (2016), galaxy protoclusters are defined as overdensities of galaxies that will eventually collapse into a galaxy cluster - i.e., a dynamically relaxed structure more massive than $10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$ at $z=0$. We adopt this definition to identify the protocluster population analyzed in this study, with the caveat that our minimum mass threshold is $10^{14.3}\leavevmode\nobreak\ {\rm M}_{\odot}$. Specifically, we define protoclusters as the ensemble of galaxies that will reside within $R_{\rm{200c}}$ at $z=0$. To ensure sufficient stellar mass resolution, we include only galaxies, regardless of their redshift, that have reached a minimum stellar mass of ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. This choice implies that the number of protocluster members will generally increase as redshift decreases, with higher-redshift protoclusters having fewer members that meet the minimum stellar mass threshold.
@@
Fig. 2 shows the spatial distribution of protocluster galaxies at different redshifts, organized by $z=0$ cluster mass: massive clusters (first and second rows), intermediate-mass clusters (third and fourth rows), and low-mass clusters (fifth and sixth rows). Black circles indicate galaxies with ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ (red circle) at $z=0$, forming our “baseline” protocluster population. Gray circles represent galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside between $R_{\rm{200c}}<R<10\leavevmode\nobreak\ R_{\rm{200c}}$ at $z=0$. From left to right, the columns show the spatial distribution of these galaxies across six snapshots from $z=0$ to $z=5$. A defining characteristic of protoclusters is their vast spatial extent, spanning several physical megaparsecs (or tens of comoving megaparsecs) at early cosmic times. This figure highlights a major observational challenge: obtaining high-quality spectroscopic redshifts over a sufficiently large area to accurately distinguish galaxies that will eventually belong to a virialized cluster at $z=0$ from those that will not. To further explore this, in Appendix A we examine the observational feasibility of the protocluster definition used in this work, specifically quantifying contamination from galaxies that may be identified as part of a protocluster but will not reside within the cluster’s virial radius at $z=0$.
@@ 2.4 Characterizing Protocluster Sizes
This definition is most similar to that employed by Muldrew et al. (2015), who defined $R_{90}$ as the radius enclosing $90\%$ of the stellar mass of their simulated protoclusters, which were sourced from the Millennium Simulation (Springel et al., 2005) with the Guo et al. (2011) semi-analytic model applied. However, unlike in Muldrew et al. (2015), where $R_{90}$ was measured with respect to the center of the galaxy cluster at $z=0$, we measure $R_{10}$ and $R_{90}$ with respect to the center of mass of the protocluster at a given redshift.
@@ 2.5 Constructing Protocluster Galaxy Density Maps
Observationally, the local environment of galaxy clusters and protoclusters has been characterized using a variety of methods that estimate galaxy density in three-dimensional space (RA, Dec, and redshift), such as nearest neighbors (e.g., Polletta et al., 2021; Champagne et al., 2021), Friends-of-Friends algorithms (e.g., Calvi et al., 2021; Helton et al., 2024a), tessellation-based density estimators (e.g., Ramella et al., 2001; Cooper et al., 2005; Cucciati et al., 2014; Darvish et al., 2015; Lemaux et al., 2018; Hung et al., 2020; Sarron & Conselice, 2021; Forrest et al., 2023), and Gaussian kernel density estimation (KDE) (e.g., Bădescu et al., 2017; McConachie et al., 2022; Brinch et al., 2023; McConachie et al., 2025). To probe the local environment of the simulated protoclusters in this study, we use adaptive binning with nearest-neighbor weighting. Specifically, we construct a three-dimensional grid using the $x$, $y$, and $z$ positions of individual galaxies from a given redshift snapshot, with a uniform bin size in comoving Mpc (cMpc) that spans the entire protocluster. The grid is constructed in cMpc, but for visualization purposes the density maps are scaled to physical Mpc (pMpc) to better capture the protocluster environment at different redshifts, given the smaller dynamic range in physical units. Moreover, while the bin size is fixed at given redshift, we adaptively adjust it to account for protocluster size evolution, setting it to $0.5$ cMpc for $10<z<$ 1, $0.75$ cMpc for $41\leq z<$ 4, and $1$ cMpc for $z\geq 4$.
@@
Using this grid, we create a multi-dimensional histogram to count galaxies in each cell and measure galaxy number density. To incorporate the local galaxy environment, we calculate the three-dimensional distance to the nearest neighbor for each galaxy and apply an exponential decay weighting, where the weight decreases with increasing neighbor distance. Specifically, the weight for each galaxy is $w=\exp(-d_{\mathrm{nn}}/\lambda_{\mathrm{nn}})$, where $\lambda_{\mathrm{nn}}$ is a parameter controlling the rate of decay, which we fix to a value of $10$ cMpc. This approach yields results quantitatively similar to the KDE approach but with significantly lower computational cost.
@@
The final density map is generated by applying a Gaussian filter to the weighted 3D histogram and summing galaxy densities along the $z$-axis to create a 2D projected density field. The highest-density region is identified as the maximum density value within this field. To locate it, we first find the maximum density in the $x$-$y$ plane and then trace along the $z$-axis to the cell with the highest galaxy concentration. Finally, the highest-density peak is defined as the centroid of the galaxies in this region.
@@
Figure 3: Top: Protocluster size evolution in physical units, characterized by $R_{90}$ and $R_{10}$, which represent the radii, measured relative to the center of mass of the protocluster at the observed redshift, that enclose $90\%$ and $10\%$, respectively, of the total stellar mass of the baseline protocluster population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$. The purple, blue, and green lines indicate the mean values in bins of $z=0$ halo mass, with bands showing the corresponding $68\%$ quantiles. Middle: Protocluster size evolution in comoving units. Bottom: Protocluster size evolution in units of arcminutes. Open markers show the angular sizes of spectroscopically-confirmed protoclusters from the literature with $>10$ members at $52<z<$ 5. Sizes correspond to the aperture radius or half the field of view; for asymmetric fields, we use half the average of both dimensions. Data and references are in provided in Table 1 of Appendix B. These results suggest that the vast majority of existing spectroscopically-confirmed protoclusters do not fully capture the expected volume occupied by the progenitors of the most massive galaxy clusters.
@@
Figure 4: The projected galaxy density distribution at $z=3$ for 15 protoclusters that will collapse into clusters by $z=0$. The density maps are constructed using adaptive binning with nearest-neighbor weighting, smoothed with a Gaussian filter, summed along the $z$-axis, and visualized using filled contours. The rows depict the projected galaxy density distribution for the progenitors of the five most massive clusters at $z=0$ (top row), five intermediate-mass clusters (middle row), and the five least massive clusters at $z=0$ (bottom row). Individual galaxies are shown as white dots, with their sizes scaling with stellar mass. Here the galaxies represent the “baseline” population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$. We highlight four regions of interest: the center of mass (red X), the highest galaxy density peak (red cross), the most massive galaxy at the given redshift (red square), and the progenitor of the $z=0$ BCG (red star). We also overplot $R_{10}$ for each protocluster, illustrating how this radial extent varies among protoclusters at this epoch. A key takeaway is that the location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) does not always correspond to the location of the most massive galaxy or the BCG progenitor.
@@
Figure 6: Three-dimensional separation between the baseline highest galaxy density peak and four regions of interest as a function of redshift: (i) the most massive galaxy at a given redshift (top-left), (ii) the BCG progenitor at a given redshift (top-right), (iii) the ${M}_{\star}$-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) highest density peak (bottom-left), and (iv) the SFR-limited (SFR $>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) highest density peak (bottom-right). The separations are binned by cluster halo mass at $z=0$, with the purple, blue, and green lines representing the median separations for progenitors of massive, intermediate, and low-mass clusters, respectively, and the error bars showing the 16th to 84th percentiles. While there is significant protocluster-to-protocluster variation in the highest density peak separation, in general, the baseline highest galaxy density peak does not coincide with the most massive galaxy or BCG progenitor at $z>1$. Likewise, at $z>2$, the highest galaxy density peaks identified by ${M}_{\star}$- and SFR-limited subpopulations generally do not trace the baseline highest galaxy density peak.
Figure 7: Cumulative recovery fraction of the baseline highest galaxy density peak as a function of the three-dimensional separation between the baseline highest galaxy density peak and the four regions of interest highlighted in Fig. 6. Solid lines show the cumulative recovery fractions in redshift bins from $z=1$ to $z=5$. In all cases at $z>2$, the recovery fractions are low, typically $\lesssim 35-40\%$ within an accuracy of 1.0 pMpc. This indicates that the most massive galaxy and BCG progenitors, as well as the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited subpopulations, do not typically coincide with the baseline highest galaxy density peak on scales smaller than 1 pMpc at $z>2$.
@@ 3.1 Comparing Sizes of Observed and Simulated Protoclusters
In the bottom panel of Fig. 3 open markers indicate the angular sizes of spectroscopically-confirmed protoclusters from the literature, with redshifts in the range $52<z<$ 5 and each containing at least $10$ spectroscopic members. These sizes correspond to either the aperture radius used to define the protocluster or half the field of view. When the field of view has unequal length and width, we use half of the average of the two dimensions. The only exception is the study by Shah et al. (2024), where the sizes are derived from the protocluster volumes provided in their work. This data, along with the corresponding references, are provided in Table 1 of Appendix B.
This comparison highlights that most spectroscopically-confirmed protoclusters in the literature do not probe the full expected volume of the progenitors of the most massive clusters at $z=0$. Instead, most of these studies are limited to aperture sizes that correspond to the protocluster core, or they lie in a region of confusion around $\sim 4\arcmin$, where they could be probing either the core of a massive cluster progenitor or the full extent of a low-mass cluster progenitor. As a result the inferences drawn from these spectroscopically-confirmed protocluster populations may be misleading or incomplete, as these surveys lack the necessary field of view and sensitivity to capture the bulk of the underlying protocluster population at these redshifts, especially for the most massive cluster progenitors.
@@
A key takeaway from Fig. 4 is that at $z=3$ the location of the highest galaxy density peak does not always correspond to the location of the most massive galaxy or the BCG progenitor. This highlights the diverse evolutionary states of protoclusters at a given redshift. Moreover, it suggests that the region of a protocluster with the highest galaxy concentration is not necessarily traced by the most massive galaxies. While the highest density peaks shown here are measured for protocluster populations with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, current observational surveys do not achieve this level of stellar mass completeness over the volumes spanned by protoclusters at $z=3$. Since current spectroscopically-confirmed protoclusters under-sample the faint end of the galaxy luminosity function, the inferred galaxy density distributions – and consequently, the highest galaxy density peaks – may not accurately reflect the underlying distribution. In other words, the true highest galaxy density region may be missed entirely by observed protocluster samples.
@@
For this cluster, we observe a discrepancy between the locations of the highest galaxy density peaks traced by the baseline population and the stellar mass-limited subpopulation at $z>2$, whereas for the SFR-limited subpopulation, the discrepancy occurs at $z=1$ and $z=5$. At $z=5$ the stellar mass- and SFR-limited subpopulations fail to recover the baseline density peak primarily because their selection functions exclude fainter and lower-mass protocluster members. However, at $z=1$, the SFR-limited subpopulation fails to recover the baseline density peak as the underlying cluster population becomes increasingly dominated by quiescent galaxies.
@@
In comparing the separation between the location of the baseline highest galaxy density peak and the location of the most massive galaxy, we find that above $z>1$ these two locations show very little overlap. The magnitude of this separation correlates modestly with the cluster halo mass at $z=0$. Similarly, the separation between the baseline highest galaxy density peak and the BCG progenitor shows little agreement beyond $z>1$, with the degree of separation again correlating with the mass of the cluster at $z=0$. In both cases, there is significant protocluster-to-protocluster variation, indicated by the 1-$\sigma$ spread in highest-density peak separations at fixed redshift, largely driven by differences in protocluster evolutionary states.
For the stellar mass-limited subpopulation (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), the separation between the highest galaxy density peak and the baseline peak exceeds $1$ pMpc at $z>2$, showing a relatively milder dependence on the cluster halo mass at $z=0$. The most striking difference occurs when comparing the baseline highest galaxy density peak to the peak probed by the SFR-limited subpopulation ($\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$). Despite sharing the same stellar mass limit (${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), we find that the overlap is poor at all redshifts. This suggests that observations focusing on highly star-forming galaxies will miss the highest concentration of galaxies during the protocluster stage of cluster assembly.
@@
In Fig. 7, we show the cumulative recovery fraction of the baseline highest galaxy density peak as a function of the three-dimensional separation from this peak to four key locations: the most massive galaxy (top-left panel), the BCG progenitor (top-right panel), and the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited populations (bottom-left and right panels, respectively). The recovery fraction is defined as the ratio of samples that recover the baseline highest galaxy density peak within a given separation to the total number of samples. As indicated by the five colored lines in Fig. 7, the results are presented as cumulative sums across five redshift snapshots from $z=1$ to $z=5$. In general, the baseline highest galaxy density peak is only recovered in $\lesssim 35-40\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2-2.6\arcmin$) at $z>2$. In other words, the most massive galaxies, BCG progenitors, and the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited populations (constrained to ${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ and SFR$>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$, respectively) do not typically coincide with the true highest galaxy density peak on scales smaller than 1 pMpc at $z>2$.
Figure 8: Three-dimensional separation as a function of redshift between the baseline highest galaxy density peak and peaks inferred from ${M}_{\star}$- and SFR-limited subsamples (left and right columns, respectively). Rows correspond to protoclusters binned by halo mass at $z=0$, with progenitors of the most (least) massive clusters shown in the top (bottom) row. The highest galaxy density peaks in the ${M}_{\star}$-limited subsamples are identified using stellar mass thresholds ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. Similarly, the SFR-limited peaks are selected using SFR thresholds from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$, with an additional stellar mass restriction of ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. While there is substantial protocluster-to-protocluster variation in peak-centric separations, as indicated by the 1-$\sigma$ error bars, both the ${M}_{\star}$- and SFR-limited subpopulations generally struggle to recover the baseline highest density peak at $z>2$, with median separations exceeding 1 pMpc even for the least restrictive thresholds.

### v2

@@ 2.3 Simulated Protocluster Population
In the review paper by Overzier (2016), galaxy protoclusters are defined as overdensities of galaxies that will eventually collapse into a galaxy cluster - i.e., a dynamically relaxed structure more massive than $10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$ at $z=0$. We adopt this definition to identify the protocluster population analyzed in this study, with the caveat that our minimum mass threshold is $10^{14.2}\leavevmode\nobreak\ {\rm M}_{\odot}$. Specifically, we define protoclusters as the ensemble of galaxies that will reside within $R_{\rm{200c}}$ at $z=0$. This includes galaxies that are no longer identifiable as distinct systems at $z=0$, having either merged with a more massive galaxy or been disrupted prior to this epoch. To ensure sufficient stellar mass resolution, we include only galaxies, regardless of their redshift, that have reached a minimum stellar mass of ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. This choice implies that the number of protocluster members will generally increase as redshift decreases, with higher-redshift protoclusters having fewer members that meet the minimum stellar mass threshold.
@@
Fig. 2 shows the projected spatial distribution of protocluster galaxies, measured relative to the protocluster’s center of mass, at different redshifts organized by $z=0$ cluster mass: massive clusters (first and second rows), intermediate-mass clusters (third and fourth rows), and low-mass clusters (fifth and sixth rows). Black circles represent galaxies with ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that are progenitors of galaxies residing within $R_{\rm{200c}}$ (red circle) at $z=0$. This includes galaxies that will merge into more massive galaxies or become disrupted prior to $z=0$, and collectively this defines our “baseline” protocluster population. Similarly, gray circles represent galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that are the progenitors of the galaxies that reside between $R_{\rm{200c}}<R<10\leavevmode\nobreak\ R_{\rm{200c}}$ at $z=0$. From left to right, the columns show the spatial distribution of these galaxies across six snapshots from $z=0$ to $z=5$. A defining characteristic of protoclusters is their vast spatial extent, spanning several physical megaparsecs (or tens of comoving megaparsecs) at early cosmic times. This figure highlights the observational challenge of obtaining high-quality spectroscopic redshifts over a sufficiently large area, as such data are required to accurately distinguish galaxies that will eventually belong to a virialized cluster at $z=0$ from those that will not.
To further explore this, in Appendix A we examine the observational feasibility of the protocluster definition used in this work, specifically quantifying contamination from galaxies that may be identified as part of a protocluster but will not reside within the cluster’s virial radius at $z=0$. The main takeaway from this comparison is that, above $z\sim 1$, the contamination fraction is moderate, increasing with center of mass centric separation and typically peaking around $\sim 15-20\%$ in the outskirts of the protocluster. This holds regardless of whether the center of mass is computed using only the baseline protocluster members (i.e., black circles in Fig. 2) or all galaxies in this region (i.e., black and gray circles in Fig. 2). Moreover, the results presented in this study are largely invariant to the adopted protocluster population definition. Specifically, we find similar results when using a more inclusive definition for the protocluster population, where members are defined as progenitors of any galaxy that is gravitationally bound to one of the 352 clusters at $z=0$, as identified by the Friends-of-Friends and Subfind algorithms mentioned in §2.1.
@@ 2.4 Characterizing Protocluster Sizes
This definition is most similar to that employed by Muldrew et al. (2015), who defined $R_{90}$ as the radius enclosing $90\%$ of the stellar mass of their simulated protoclusters, which were sourced from the Millennium Simulation (Springel et al., 2005) with the Guo et al. (2011) semi-analytic model applied. However, unlike in Muldrew et al. (2015), where $R_{90}$ was measured with respect to the center of the galaxy cluster at $z=0$, we measure $R_{10}$ and $R_{90}$ with respect to the center of mass of the ensemble of the protocluster members a given redshift.
@@ 2.5 Constructing Protocluster Galaxy Density Maps
Observationally, the local environment of galaxy clusters and protoclusters has been characterized using a variety of methods that estimate galaxy density in three-dimensional space (RA, Dec, and redshift), such as nearest neighbors (e.g., Polletta et al., 2021; Champagne et al., 2021), Friends-of-Friends algorithms (e.g., Calvi et al., 2021; Helton et al., 2024a), tessellation-based density estimators (e.g., Ramella et al., 2001; Cooper et al., 2005; Cucciati et al., 2014; Darvish et al., 2015; Lemaux et al., 2018; Hung et al., 2020; Sarron & Conselice, 2021; Forrest et al., 2023), and Gaussian kernel density estimation (KDE) (e.g., Bădescu et al., 2017; McConachie et al., 2022; Brinch et al., 2023; McConachie et al., 2025). To probe the local environment of the simulated protoclusters in this study, we use adaptive binning with nearest-neighbor weighting, chosen to approximate the aforementioned observational approaches for quantifying protocluster galaxy density fields. Specifically, we construct a three-dimensional grid using the $x$, $y$, and $z$ positions of individual galaxies from a given redshift snapshot, with a uniform bin size in physical Mpc (pMpc) that spans the entire protocluster. While the bin size is fixed at a given redshift, we adaptively adjust it across redshifts to account for protocluster size evolution, setting it to one-fifth of the median value of $R_{10}(z,M_{\rm{200}}^{z=0})$, which corresponds to bin sizes ranging from $\sim 0.1-0.2$ pMpc.
@@
Using this grid, we create a multi-dimensional histogram to count galaxies in each cell and measure galaxy number density. To incorporate the local galaxy environment, we calculate the three-dimensional distance to the nearest neighbor for each galaxy and apply an exponential decay weighting, where the weight decreases with increasing neighbor distance. Specifically, the weight for each galaxy is $w=\exp(-d_{\mathrm{nn}}/\lambda_{\mathrm{nn}})$, where $\lambda_{\mathrm{nn}}$ is a parameter controlling the rate of decay, which we fix to a value of $1$ pMpc. This approach yields results quantitatively similar to the KDE approach but with significantly lower computational cost.
@@
The final density map is generated by applying a Gaussian filter to the weighted 3D histogram and summing galaxy densities along the $z$-axis to create a 2D projected density field. This approach approximates observational methods, which estimate projected galaxy densities by summing over galaxies in redshift slices. The highest-density region is identified as the maximum density value within this field. To locate it, we first find the maximum density in the $x$-$y$ plane and then trace along the $z$-axis to the cell with the highest galaxy concentration. Finally, the highest-density peak is defined as the centroid of the galaxies in this region. When comparing the 2D (projected) highest galaxy density peak with the 3D maximum density peak – defined as the bin in the 3D density field with the highest number density – we find that across the redshift range $z=2$ to $z=5$ that the median spatial offset between the two peaks ranges from $0.20$ and $0.30$ pMpc, which corresponds to $\sim 2-3$ times the minimum bin size. Nevertheless, we adopt the projected density peaks in our analysis, as it more closely resembles how density peaks are identified in observations.
@@
Figure 3: Top: Protocluster size evolution in physical units, characterized by $R_{90}$ and $R_{10}$, which represent the radii, measured relative to the center of mass of the ensemble of protocluster members at the observed redshift, that enclose $90\%$ and $10\%$, respectively, of the total stellar mass of the baseline protocluster population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$. The purple, blue, and green lines indicate the mean values in bins of $z=0$ halo mass, with bands showing the corresponding $68\%$ quantiles. Middle: Protocluster size evolution in comoving units. Bottom: Protocluster size evolution in units of arcminutes. Open markers show the angular sizes of spectroscopically confirmed protoclusters from the literature with $>10$ members at $52<z<$ 5. Sizes correspond to the aperture radius or half the field of view; for asymmetric fields, we use half the average of both dimensions. Data and references are in provided in Table 1 of Appendix B. These results suggest that the vast majority of existing spectroscopically confirmed protoclusters do not fully capture the expected volume occupied by the progenitors of the most massive galaxy clusters.
@@
Figure 4: The projected galaxy density distribution at $z=3$ for 15 protoclusters that will collapse into clusters by $z=0$. The density maps are constructed using adaptive binning with nearest-neighbor weighting, smoothed with a Gaussian filter, summed along the $z$-axis, and visualized using filled contours. The rows depict the projected galaxy density distribution for the progenitors of the five most massive clusters at $z=0$ (top row), five intermediate-mass clusters (middle row), and the five least massive clusters at $z=0$ (bottom row). Individual galaxies are shown as white dots, with their sizes scaling with stellar mass. Here the galaxies represent the “baseline” population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that are the progenitors of the galaxies that reside within $R_{\rm{200c}}$ at $z=0$. We highlight four regions of interest: the center of mass (red X), the highest galaxy density peak (red cross), the most massive galaxy at the given redshift (red square), and the progenitor of the $z=0$ BCG (red star). We also overplot $R_{10}$ for each protocluster, illustrating how this radial extent varies among protoclusters at this epoch. A key takeaway is that the location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) does not always correspond to the location of the most massive galaxy or the BCG progenitor.
@@ 3.1 Comparing Sizes of Observed and Simulated Protoclusters
The open markers shown in the bottom panel of Fig. 3 indicate the angular sizes of spectroscopically confirmed protoclusters from the literature, with redshifts in the range $52<z<$ 5 and each containing at least $10$ spectroscopic members. These sizes correspond to either the aperture radius used to define the protocluster or half the field of view. When the field of view has unequal length and width, we use half of the average of the two dimensions. The only exception is the study by Shah et al. (2024), where the sizes are derived from the protocluster volumes provided in their work. This data, along with the corresponding references, are provided in Table 1 of Appendix B.
This comparison highlights that most spectroscopically confirmed protoclusters in the literature do not probe the full expected volume of the progenitors of the most massive clusters at $z=0$. Instead, a large fraction of these studies are limited to radial aperture sizes less than $4\arcmin$, which are insufficient for capturing beyond the innermost region of even the progenitors of the least massive clusters at $z=0$. As a result the inferences drawn from these spectroscopically confirmed protocluster populations may be misleading or incomplete, as these surveys lack the necessary field of view and sensitivity to capture the bulk of the underlying protocluster population at these redshifts, especially for the most massive cluster progenitors. Moreover, as shown in Chiang et al. (2017), in order to probe the full contribution of protoclusters to the cosmic star formation rate density, and also to driving reionization, the fully extent of the protocluster, as opposed to the core, is required. Lim et al. (2024) also showed that estimates of $z=0$ halo masses are sensitive to the aperture size used to define the protocluster.
@@
A key takeaway from Fig. 4 is that at $z=3$ the location of the highest galaxy density peak does not always correspond to the location of the most massive galaxy or the BCG progenitor. This highlights the diverse evolutionary states of protoclusters at a given redshift. Moreover, it suggests that the region of a protocluster with the highest galaxy concentration is not necessarily traced by the most massive galaxies. While the highest density peaks shown here are measured for protocluster populations with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, current observational surveys do not achieve this level of stellar mass completeness over the volumes spanned by protoclusters at $z=3$. Since current spectroscopically confirmed protoclusters under-sample the faint end of the galaxy luminosity function, the inferred galaxy density distributions – and consequently, the highest galaxy density peaks – may not accurately reflect the underlying distribution. In other words, the true highest galaxy density region may be missed entirely by observed protocluster samples. This could have important implications for how galaxy properties – e.g., stellar mass and star formation rate – vary with local galaxy density in protocluster environments, which are trends that have been explored in observations (e.g., Shimakawa et al., 2018; Koyama et al., 2021).
@@
For this cluster, we observe a discrepancy between the locations of the highest galaxy density peaks traced by the baseline population and the stellar mass-limited subpopulation at $z=3$, whereas for the SFR-limited subpopulation, the discrepancy only occurs at $z=1$. The former can be attributed to the selection function excluding low-mass protocluster members at $z=3$, whereas the latter reflects that at $z=1$ the densest regions are dominated by quiescent galaxies. Nevertheless, for this specific case, excluding low-mass and faint protocluster members has a minor impact on recovering the baseline highest density peak.
Figure 6: Three-dimensional separation between the baseline highest galaxy density peak and four regions of interest as a function of redshift: (i) the most massive galaxy at a given redshift (top-left), (ii) the BCG progenitor at a given redshift (top-right), (iii) the ${M}_{\star}$-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) highest density peak (bottom-left), and (iv) the SFR-limited (SFR $>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) highest density peak (bottom-right). The separations are binned by cluster halo mass at $z=0$, with the purple, blue, and green lines representing the median separations for progenitors of massive, intermediate, and low-mass clusters, respectively, and the error bars showing the 16th to 84th percentiles. While there is significant protocluster-to-protocluster variation in the highest density peak separation, in general, the baseline highest galaxy density peak deviates from the most massive galaxy and BCG progenitor at $z>1$, with median separations exceeding 1.0 pMpc. The highest galaxy density peaks identified by ${M}_{\star}$-limited (SFR-limited) subpopulations exhibit median separations from the baseline highest galaxy density peak that increase (decrease) with increasing redshift.
@@
[…]
For the stellar mass-limited subpopulation (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), the median separation between the highest galaxy density peak and the baseline peak increases from $0.25$ to $1.0$ pMpc between $z=2$ and $z=5$, and exhibits a relatively milder dependence on the cluster halo mass at $z=0$. The opposite trend is observed when comparing with the SFR-limited subpopulation ($\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$), where the median separation decreases from $1.0$ to $0.5$ pMpc between $z=2$ and $z=5$.
[…]
@@
In Fig. 7, we show the cumulative recovery fraction of the baseline highest galaxy density peak as a function of the three-dimensional separation from this peak to four key locations: the most massive galaxy (top-left panel), the BCG progenitor (top-right panel), and the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited populations (bottom-left and right panels, respectively). The recovery fraction is defined as the ratio of samples that recover the baseline highest galaxy density peak within a given separation to the total number of samples.
As indicated by the five colored lines in Fig. 7, the results are presented as cumulative sums across five redshift snapshots from $z=1$ to $z=5$. We find that the baseline highest density peak coincides with the most massive galaxy and BCG progenitors in $\lesssim 45-55\%$ of the cases within an accuracy of 1.0 pMpc ($\sim 2-2.6\arcmin$) at $z>2$. This agreement is improved when measured relative to the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited population, which recover the baseline highest density peak in $\lesssim 55-65\%$ of the cases within an accuracy of 1.0 pMpc at $z>2$. In other words, on scales smaller than $1.0$ pMpc at $z>2$ – which is approximately the median value of $R_{10}$ at $z>2$ – the most massive galaxy and BCG progenitor are associated with the highest density peak about half the time, and relative to observationally-limited subpopulations, more than half of the peaks coincide with the true highest galaxy density peak.
[…]